OpenStudy (acespeedfighter):

What is the value of the expression 6^2 + 3^3? (2 points)

OpenStudy (acespeedfighter):

i think its 63 but im not too sure

OpenStudy (welshfella):

6^2 = 6 * 6 3^3 = 3 * 3* 3

OpenStudy (acespeedfighter):

so add them uo right?

OpenStudy (acespeedfighter):

@welshfella

OpenStudy (welshfella):

yes 36 + 27 = 63

OpenStudy (acespeedfighter):

oh ok thx for checking

OpenStudy (welshfella):

yw